# Contour Flow Constraint: Preserving Global Shape Similarity for Deep Learning based Image Segmentation

### v1

@@ 6.3 Validate Effectiveness of the Shape Loss
To validate the effectiveness of the proposed shape loss in preserving global shape similarity between the network’s segmentation and the ground truth, we design the following experiments. First, we train the benchmark segmentation frameworks with two baseline loss functions, i.e. cross-entropy loss $\mathcal{L}_{CE}$ and dice loss $\mathcal{L}_{Dice}$. Then, we train the networks once again with the same setups but adding the shape loss $\mathcal{L}_{S}$ to baseline losses. Finally, we evaluate and compare their performances on test datasets by the three metrics, namely IOU, BD, BDSD. By the comparison, we shall see whether networks trained with the shape loss exhibit better global shape similarity and whether such an effect depends on specific network architecture.
@@ 6.3.3 Performances and Analyses
Table 1 presents the segmentation performance results of the benchmark networks on the test sets of various datasets. With incorporation of the proposed shape loss, the network’s performance significantly improves on all datasets with few exceptions. Figures 6 provides a more intuitive visualization, showing that on the same test image, the networks trained with the shape loss produce segmentation results that are much closer to the ground truth in terms of shape. Moreover, these results exhibit fewer salt-and-pepper noise, isolated points, and erroneous regions. This indicates that training with the shape loss allows the network parameters to capture the global shape information of objects, resulting in improved segmentation performance.
These comparative experiments also demonstrate that the proposed shape loss is not dependent on specific network architectures or baseline loss functions when applied in practice. As long as this loss term is incorporated into the network training process, segmentation performance can be improved. In essence, it can be seamlessly integrated into arbitrary networks’ training progress, enhancing segmentation performance across the board. This showcases the proposed shape loss’s generalizability.
@@ 6.4.3 Performances and Analyses
Table 2 presents the test results on various datasets. It can be observed that the proposed CFSSnet exhibits greater robustness to noise. This is because CFSSnet enforces global shape similarity constraints through its network architecture, which is derived from the iterative schemes of the variational model through unrolling. Consequently, during the forward propagation, the network essentially performs the iteration schemes, so that the output results inherently possess the properties of the shape similarity constraint added to the variational model.
In contrast to training with shape loss, where the networks learn the correct shape information and perform well on the original test set, they struggle to adapt to the new image distribution introduced by noise. The experiments prove that due to its inherent shape constraint ability, CFSSnet can adapt to new image distributions even in the presence of noise, and capture shape features accurately. Therefore, CFSSnet maintains the property of global shape similarity in its output segmentations. Figure 6 provides a visual comparison and demonstration of the experimental results.
@@ 6.5.1 Reference Methods
We compare our methods, i.e. $\mathcal{L}_{S}$ shape loss and CFSSnet with the most related SOTA methods, including Dcan(Chen et al., 2016), Dmtn(Tan et al., 2018), Psi-net(Murugesan et al., 2019a), Conv-MCD(Murugesan et al., 2019b), $\mathcal{L}_{perim}$ perimeter loss(Jurdi et al., 2021), $\mathcal{L}_{AB}$ active boundary loss(Wang et al., 2022). These methods all considered incorporating contours, distance maps, or boundary information as shape prior for better segmentation, and have demonstrated competitive performance on similar image segmentation tasks. To ensure a fair comparison, we adopt the same training and testing protocols across all methods. U-net(Ronneberger et al., 2015) is chosen as a baseline model for methods of adding losses. The models are initialized by trained with 150 epochs, batch sizes of 16, and a learning rate of $41e-$ 4. Then, all methods are evaluated on testset by the aforementioned metrics IOU, BD, BDSD.
@@ 6.5.2 Performances and Analyses
The results of comparisons are summarized in Table 3 Some examples are also displayed in Figure 7 to demonstrate the advantages of our proposed methods. As shown, our methods consistently outperform the compared methods across all metrics, demonstrating superior overall segmentation accuracy, better boundary alignment and precision, and a higher degree of shape preservation. These results validate the effectiveness of our approaches, particularly for scenarios that require high shape fidelity and precision.
@@
6.6 Ablation Study for CFSSnet
In this section, an ablation study is conducted to understand the individual contribution of modules in CFSSnet, i.e. the contour flow branch and shape constraint module. By doing so, we can get a better understanding of where the inherent shape-preserving ability of CFSSnet comes from.
@@ 6.6 Ablation Study for CFSSnet
To proceed, we modify $T$, the number of shape constraints modules from no such modules at all, and increase $10$ modules each time until $T=40$ to check the best choice of module number to implement. For a contribution of the contour flow branch, we do comparable experiments to see how metrics vary with and without the branch. If there is no such branch, $\boldsymbol{F}=[\cos(\textrm{conv}(U)),\sin(\textrm{conv}(U))]$, i.e. directly from backbone logits output. These different settings are trained and evaluated on the WBC2 dataset to see whether the inherent shape-preserving ability of CFSSnet is retained or not.
6.6.1 Why contour flow prediction branch is necessary?
@@
As shown in Table 4, the inclusion of the contour flow branch leads to a great improvement in the overall performance of CFSSnet. Without this branch, the network lacks the representational capacity to accurately predict the contour flow $\boldsymbol{F}$. Notably, even in the absence of shape constraint modules, the network benefits from learning the ground truth flow, which enhances shape similarity.
6.6.2 What is the optimum number of shape constraint modules?
@@
Our ablation study suggests that using $T=20$ shape constraint modules offers the best overall performance, which is the setting for this paper’s implementation. With this configuration, the network strikes an ideal balance between IOU, shape properties (BD, BDSD), and resistance to noise. The reasoning is that our shape constraint modules are essentially an unrolled iterative process for solving the variational segmentation model, which requires a certain number of iterations to reach an optimal solution. When $T<20$, the iterations may be insufficient. However, when $T>20$, further iterations do not provide any additional benefit. Moreover, adding unnecessary modules can overcomplicate the gradient graph during backpropagation, making it harder for network convergence. We recommend using $T$ in the range of [15, 25].

### v2

@@ 6.3 Validate Effectiveness of the Shape Loss
To validate the effectiveness of the proposed shape loss in preserving global shape similarity between the network’s segmentation and the ground truth, we design the following experiments. First, we train the benchmark segmentation frameworks with two baseline loss functions, i.e. cross-entropy loss $\mathcal{L}_{CE}$ and dice loss $\mathcal{L}_{Dice}$ (soft dice loss without squared denominator as defined in Eq.(7)). Then, we train the networks once again with the same setups but adding the shape loss $\mathcal{L}_{S}$ to baseline losses. Finally, we evaluate and compare their performances on test datasets by the three metrics, namely Dice, BD, BDSD. By the comparison, we shall see whether networks trained with the shape loss exhibit better global shape similarity and whether such an effect depends on specific network architecture.
@@ 6.3.3 Performances and Analyses
Table 1 presents the segmentation performance results of the benchmark networks on the test sets of various datasets. With incorporation of the proposed shape loss, the network’s performance significantly improves on all datasets with few exceptions. Figures 6 provides a more intuitive visualization, showing that on the same test image, the networks trained with the shape loss produce segmentation results that are much closer to the ground truth in terms of shape. Moreover, these results exhibit fewer salt-and-pepper noise, isolated points, and erroneous regions. This indicates that training with the shape loss allows the network parameters to capture the global shape information of objects, resulting in improved segmentation performance. These comparative experiments also demonstrate that the proposed shape loss is not dependent on specific network architectures or baseline loss functions when applied in practice. As long as this loss term is incorporated into the network training process, segmentation performance can be improved. In essence, it can be seamlessly integrated into arbitrary networks’ training progress, enhancing segmentation performance across the board, which showcases the proposed shape loss’s generalizability.
@@ 6.4.3 Performances and Analyses
Table 2 presents the test results on various datasets. It can be observed that the proposed CFSSnet exhibits greater robustness to noise. This is because CFSSnet enforces global shape similarity constraints through its network architecture, which is derived from the iterative schemes of the variational model through unrolling. Consequently, during the forward propagation, the network essentially performs the iteration schemes, so that the output results inherently possess the properties of the shape similarity constraint added to the variational model. In contrast to training with shape loss, where the networks learn the correct shape information and perform well on the original test set, they struggle to adapt to the new image distribution with noise. Figure 6 provides a visual comparison and demonstration of the experimental results.
@@
6.5 3D Segmentation with the Proposed Method.
In the field of medical image processing, 3D segmentation is commonly used to handle the detailed structures in medical images. We find that, by using an equivalent constraint $\nabla u\times\nabla\phi_{g}=\boldsymbol{0}$, our proposal can also implemented in 3D. For some preliminary results, We carried out 3D segmentation experiments on PROMISE2012 challenge dataset (Litjens et al., 2014) with two baseline networks, 3D-Unet (Çiçek et al., 2016) and V-net (Milletari et al., 2016). Likewise in section 6.3, for each network, three types of loss, namely the weighted CE loss $\mathcal{L}_{WCE}$, dice loss $\mathcal{L}_{Dice}$, and Tversky loss (Salehi et al., 2017) $\mathcal{L}_{Tversky}$ are applied and then compared with adding our 3D shape loss $\mathcal{L}_{3DS}=\|\nabla u\times\nabla\phi_{g}\|$. The results are shown in Table 3. Refer to Figure 7 for intuitions of our method’s advantages in maintaining boundary shapes and eliminating outer isolated regions or inner holes.
Table 3: Preliminary results of 3D Shape Loss on baseline models
Network Loss Function PROMISE2012
[…]
3D-Unet $\mathcal{L}_{WCE}$ 66.49 6.120 4.940
$\mathcal{L}_{WCE}+\mathcal{L}_{3DS}$ 68.61 2.982 2.185
$\mathcal{L}_{Dice}$ 74.59 8.897 11.25
$\mathcal{L}_{Dice}+\mathcal{L}_{3DS}$ 79.51 2.235 2.381
$\mathcal{L}_{Tversky}$ 78.39 6.903 9.295
$\mathcal{L}_{Tversky}+\mathcal{L}_{3DS}$ 82.21 1.641 1.437
V-Unet $\mathcal{L}_{WCE}$ 69.37 4.579 4.810
$\mathcal{L}_{WCE}+\mathcal{L}_{3DS}$ 77.39 2.211 1.736
$\mathcal{L}_{Dice}$ 65.02 3.245 5.285
$\mathcal{L}_{Dice}+\mathcal{L}_{3DS}$ 80.82 1.744 1.501
$\mathcal{L}_{Tversky}$ 77.68 3.938 4.852
$\mathcal{L}_{Tversky}+\mathcal{L}_{3DS}$ 81.53 1.785 1.731
Figure 7: Comparison between 3D segmentation (ground truth colored in white) on prostate MR images trained with Tversky loss alone (colored in red) and together with 3D shape loss (colored in blue).
@@ 6.6.1 Reference Methods
We compare our methods, i.e. $\mathcal{L}_{S}$ shape loss and CFSSnet with the most related SOTA methods, including Dcan (Chen et al., 2016), Dmtn (Tan et al., 2018), Psi-net (Murugesan et al., 2019a), Conv-MCD (Murugesan et al., 2019b), $\mathcal{L}_{perim}$ perimeter loss (Jurdi et al., 2021), $\mathcal{L}_{AB}$ active boundary loss (Wang et al., 2022). These methods all considered incorporating contours, distance maps, or boundary information as shape prior for better segmentation, and have demonstrated competitive performance on similar image segmentation tasks. To ensure a fair comparison, we adopt the same training and testing protocols across all methods. U-net (Ronneberger et al., 2015) is chosen as a baseline model for methods of adding losses. The models are initialized by trained with 150 epochs, batch sizes of 16, and a learning rate of $41e-$ 4. Then, all methods are evaluated on testset by the aforementioned metrics Dice, BD, BDSD.
@@ 6.6.2 Performances and Analyses
The results of comparisons are summarized in Table 4. Some examples are also displayed in Figure 8 to demonstrate the advantages of our proposed methods. As shown, our methods consistently outperform the compared methods across all metrics, demonstrating superior overall segmentation accuracy, better boundary alignment and precision, and a higher degree of shape preservation.
@@
6.7 Summarized Analysis
An important observation drawn from experiments is that the proposed shape loss $\mathcal{L}_{S}$ and corresponding variational iterations (Algorithm 1) have the main impact and primarily drives improvements in boundary-distance metrics (BD, BDSD) rather than overlap-based Dice. Such improvements are consistent across datasets, suggesting that the proposed shape constraint consistently enforces tighter adherence to the true contour across different segmentation tasks and imaging modalities. The reason for this discrepancy is that our shape constraint is derived from definition of global shape similarity and explicitly targets boundary alignment, so it directly penalizes deviations along boundary. By contrast, Dice depends on the overall region-overlap, which is indirectly improved by including the proposed shape constraint.
6.8 Ablation Study
6.8.1 Comparison for Shape Loss with Established Loss
As demonstrated in (Isensee et al., 2021), the most common loss function for medical segmentation is the sum of cross-entropy and Dice loss (DiceCE). Although in previous sections we have conducted extensive experiments on adding proposed shape loss onto cross-entropy loss and Dice loss alone, we are still interested in comparison of shape loss with this established loss function. We use U-net as baseline, and compare several different combination of loss functions on the BUSI and Kvasir datasets in Table 5. One can observe dice loss in general yields better Dice score, but by adding shape loss, shape metrics can still have improvement.
Table 5: Comparison on different combinations of shape loss and established loss
Loss BUSI Kvasir
[…]
$\mathcal{L}_{DiceCE}$ 75.18 13.47 10.74 76.46 13.06 11.89
$\mathcal{L}_{CE}+\mathcal{L}_{S}$ 75.76 12.58 9.740 74.54 13.73 12.35
$\mathcal{L}_{Dice}+\mathcal{L}_{S}$ 76.18 13.48 10.84 76.83 12.86 11.67
$\mathcal{L}_{DiceCE}+\mathcal{L}_{S}$ 76.69 13.04 9.486 76.35 12.26 11.20
6.8.2 Contour Flow Branch and Shape Constraint Module in CFSSnet
@@
In this section, an ablation study is conducted to understand the individual contribution of modules in CFSSnet, i.e. the contour flow branch and shape constraint module. By doing so, we can get a better understanding of where the inherent shape-preserving ability of CFSSnet comes from. To proceed, we modify $T$, the number of shape constraints modules from no such modules at all, and increase $10$ modules each time until $T=40$ to check the best choice of module number to implement. For a contribution of the contour flow branch, we do comparable experiments to see how metrics vary with and without the branch. If there is no such branch, $\boldsymbol{F}=[\cos(\textrm{conv}(U)),\sin(\textrm{conv}(U))]$, i.e. directly from backbone logits output. These different settings are trained and evaluated on the WBC2 dataset to see whether the inherent shape-preserving ability of CFSSnet is retained or not.
@@
Why contour flow prediction branch is necessary? As shown in Table 6, the inclusion of the contour flow branch leads to a great improvement in the overall performance of CFSSnet. Without this branch, the network lacks the representational capacity to accurately predict the contour flow $\boldsymbol{F}$. Notably, even in the absence of shape constraint modules, the network benefits from learning the ground truth flow, which enhances shape similarity.
@@
What is the optimum number of shape constraint modules? Our ablation study suggests that using $T=20$ shape constraint modules offers the best overall performance, which is the setting for this paper’s implementation. With this configuration, the network strikes an ideal balance between Dice, shape properties (BD, BDSD), and resistance to noise. The reasoning is that our shape constraint modules are essentially an unrolled iterative process for solving the variational segmentation model, which requires a certain number of iterations to reach an optimal solution. When $T<20$, the iterations may be insufficient. However, when $T>20$, further iterations do not provide any additional benefit. Moreover, adding unnecessary modules can overcomplicate the gradient graph during backpropagation, making it harder for network convergence. We recommend using $T$ in the range of [15, 25].
@@
6.8.3 Impact of Contour Flow Estimation Errors
[…]
Figure 9: Mask Refinements by Algorithm 1 with disturbed flow.
Next, we evaluate the magnitude of flow estimation errors on the Kvasir datasets on both training and inference phases and compare two configurations: (1) our standard CF branch, equal depth to the segmentation branch, and (2) an enlarged CF branch with double depth. For each test image, we compute the predicted flow $\boldsymbol{F}$ and compare it to the ground truth flow $\boldsymbol{F}_{g}$ using three metrics: Average Cosine Similarity (ACS), End-Point Error (EPE), and Average Divergence Error (ADE). Table 7 summarizes these metrics on both phases for the two CF-branch configurations. We observe that while some error in learning $\boldsymbol{F}_{g}$ is unavoidable, current CF-branch depth has moderate estimation errors in $\boldsymbol{F}_{g}$. Using deeper CF branch has marginal improvement, but will add cost of computations.
Table 7: Contour Flow Learning Errors on Kvasir Dataset
Depth Training Inference
[…]
Equal 0.9701 0.1525 0.0391 0.8975 0.2885 0.0436
Double 0.9621 0.1801 0.0384 0.9050 0.2831 0.0421
@@ 7 Conclusion and Discussion
Despite the demonstrated strengths of our shape loss and CFSSnet framework, several limitations remain. First, incorporating the shape loss introduces additional computational overhead during both training and inference, which may hinder scalability on very large volumes or in resource-constrained settings. Second, the method could be sensitive to class imbalance: rare or small structures may not receive sufficient emphasis in the shape prior, leading to suboptimal segmentation for imbalanced classes. Third, while our approach effectively enforces global shape similarity, it does not theoretically guarantee the preservation of complex topological properties such as connectivity or genus. In future work, we plan to investigate more efficient optimization strategies to reduce computation time, develop adaptive weighting schemes to mitigate class imbalance, and integrate explicit topological constraints or priors to ensure connectivity and other critical topological invariants.
Although we have only discussed the preservation of binary classification object shapes, the same idea can be applied to multi-classification. Multi-class segmentation tasks can always be decomposed into some binary sub-problems regarding each class to implement our method. One can also extend the proposed contour flow with a softmax manner to be class-wise constrained.
Supplementary Material
Proof of Corollary 1
Proof.
$\partial\Omega_{=}\{i\in\Omega|u(i)=\gamma\in\mathbbm{R}\}$, and $\delta_{u}^{\gamma,m}$ is the $m$-th $\gamma$-contour branch of $u$. Since $u$ has contour similarity with $\boldsymbol{g}$, by definition, $\exists\delta_{g}^{\alpha,n}$ such that $\delta_{u}^{\gamma,m}=\delta_{g}^{\alpha,n}$. By definition of $\boldsymbol{g}$’s contours, $\forall i,j\in\delta_{u}^{\gamma,m},\phi_{g}(i)=\phi_{g}(j)=\alpha$. Therefore, $D(i,\partial\Omega_{g})=|\phi_{g}(i)|=|\alpha|=|\phi_{g}(j)|=D(j,\partial\Omega_{g}).$
Proof of Lemma 1
For $f\in C^{1}(\Omega)$, consider $\nabla f(i)\in\mathbbm{R}^{2}$ and $\alpha$-contour $\delta_{f}^{\alpha}=\{j\in\Omega|f(j)=\alpha\}$, where $\alpha=f(i)$. If $\delta_{f}^{\alpha}$ is complex-connected curve with several connected components, we can simply focus on the one branch where $i$ is located. Therefore, we assume $\delta_{f}^{\alpha}$ is a $C^{1}$ Jordan curve, which can be represented by $f(j)\equiv\alpha$. Introduce an arc-length parameter $t\in\mathbbm{R}$ to represent it as a parametric curve $\varphi(t)=j\in\Omega$. Then, by derivative of $t$ at both sides of $f(j)\equiv\alpha$:
$\frac{d}{dt}f(j)=\frac{d}{dt}f(\varphi(t))=\nabla f(\varphi(t))\cdot\varphi^{\prime}(t)\equiv 0=\frac{d}{dt}\alpha,$ (20)
where $\nabla f(\varphi(t))=\nabla f(j)$ is the gradient vector of $f$ at $j$, and $\varphi^{\prime}(t)$ is the tangent vector of contour branch $\delta_{f}^{\alpha}$ at $j$. This equation indicates the two are orthogonal with each other at any locations on the $\alpha$-contour, i.e. $\nabla f(i)\perp\delta_{f}^{\alpha}$.
Proof of Theorem 1
Recall contour similarity $\delta_{u}\sim\delta_{g}$ means $\forall\delta_{u}^{\beta,m}$, i.e. arbitrary contour branch of $u$, there exists $\delta_{g}^{\alpha,n}$, i.e. contour branch of $\boldsymbol{g}$, such that $\delta_{u}^{\beta,m}=\delta_{g}^{\alpha,n}$. Since $\delta_{u}^{\beta,m},\delta_{g}^{\alpha,n}$ are $C^{1}$ Jordan curve, we can always introduce an arc-length parameter $t\in\mathbbm{R}$ to represent a contour branch by parametric curve $\varphi(t)=i,\mathbbm{R}\mapsto\delta_{g}^{\alpha,n}$, where the arc-length parameter $t$ increases along the tangent direction $\boldsymbol{F}_{g}(i)=\varphi^{\prime}(t)$.
($\Longrightarrow$:) If $\delta_{u}\sim\delta_{g}$, then $\forall i\in\Omega$, denote $\beta=u(i)$, $i\in\delta_{u}^{\beta,m}$ (some branch of $\delta_{u}^{\beta}$). By definition of contour similarity, $\exists\delta_{g}^{\alpha,n}$, such that $\delta_{u}^{\beta,m}=\delta_{g}^{\alpha,n}$. Therefore, $u(i)\equiv\beta,i\in\delta_{g}^{\alpha,n}$. Now, use the parametric form to represent $\delta_{g}^{\alpha,n}$ by $\varphi(t)=i\in\delta_{g}^{\alpha,n}$. By derivative of $t$ at both sides of $u(i)\equiv\beta$:
$\frac{d}{dt}u(i)=\frac{d}{dt}u(\varphi(t))=\nabla u(\varphi(t))\cdot\varphi^{\prime}(t)=0=\frac{d}{dt}\beta,$ (21)
where $\nabla u(\varphi(t))=\nabla u(i),\varphi^{\prime}(t)=\boldsymbol{F}_{g}(i)$. To conclude, $\forall i\in\Omega,\nabla u(i)\cdot\boldsymbol{F}_{g}(i)=0$.
[…]
We denote $i^{*}=\psi(t^{*})\in\delta_{u}^{\beta^{*},m^{*}}$. Then, $H^{\prime}(t^{*})=\nabla\phi_{g}(i^{*})\cdot\psi^{\prime}(t^{*})\neq 0$. Here, $\psi^{\prime}(t^{*})$ is the tangent vector of $\delta_{u}^{\beta^{*},m^{*}}$ at $i^{*}$. According to Lemma 1, $\psi^{\prime}(t^{*})$ is orthogonal with $\nabla u(i^{*})$. Moreover, $\nabla\phi_{g}(i^{*})$ is also orthogonal with $\boldsymbol{F}_{g}(i^{*})$ by Definition 4. Thus, by rotating vectors $\nabla\phi_{g}(i^{*}),\psi^{\prime}(t^{*})$ with $\pi/2$, we get $\nabla u(i^{*})\cdot\boldsymbol{F}_{g}(i^{*})\neq 0$, which is contradicted with $\forall i\in\Omega,\nabla u(i)\cdot\boldsymbol{F}_{g}(i)=0$.
Proof of Theorem 2
Since
.$\begin{array}[]{rl}\displaystyle\biggl{.}\int_{\Omega}u~{}{\rm div}(q\boldsymbol{F}_{g})di&=-\displaystyle\biggl{.}\int_{\Omega}q(\nabla u\cdot\boldsymbol{F}_{g})di+\displaystyle\biggl{.}\int_{\partial\Omega}uq\boldsymbol{F}_{g}\cdot\vec{\boldsymbol{n}}dS\\ &=-\displaystyle\biggl{.}\int_{\Omega}q(\nabla u\cdot\boldsymbol{F}_{g})di,\end{array}$
then .$\biggl{.}\max_{q\in C^{1}(\Omega)}\left\{\int_{\Omega}u~{}{\rm div}(q\boldsymbol{F}_{g})di\right\}=0$ if $u\in\mathbb{C}$. Else, taking $q=k(\nabla u\cdot\boldsymbol{F}_{g})\in C^{1}(\Omega)$ and let $k$ go to +$\infty$+ ∞, then one can obtain .$\biggl{.}\max_{q\in C^{1}(\Omega)}\left\{\int_{\Omega}u~{}{\rm div}(q\boldsymbol{F}_{g})di\right\}=+\infty.$ which completes the proof.